OpenStudy (anonymous):

change 144 in^2 to ft^2

OpenStudy (compphysgeek):

\(144 in^2 = (12 in)^2 \) agreed? \(12 in = 1 ft\). so \(144 in^2 = 1 ft^2\)

OpenStudy (anonymous):

@compphysgeek i dnt understand :(

OpenStudy (compphysgeek):

144 in^2 = 12 in * 12 in 12 in = 1 ft 144 in^2 = 12 in * 12 in = 1 ft * 1 ft = 1 ft^2 hope that helps :) it's again all about the conversion factor

OpenStudy (anonymous):

my weaknesses s conversion :(

OpenStudy (anonymous):

can you help me

OpenStudy (compphysgeek):

it's like I said earlier. on the left hand side you start with the number that you've got. let's say \( 500\,m\) again. If you want to know how many feet that is then you multiply the number that you've got with the conversion factor \[ 500\, m \times \frac{1\, in}{0.0254\, m} \times \frac{1\, ft}{12\, in} = ?\, ft\] in the first step the m cancel out, in the second step the in cancel out. on the left side you start with m and on the right side you end up with ft

OpenStudy (anonymous):

how about this change 1ft^3 to ft^3

OpenStudy (compphysgeek):

1 ft^2 to in^3? or m^3?

OpenStudy (anonymous):



OpenStudy (compphysgeek):

ft^3 to ft^3 .. that's the same unit. there is nothing to change

OpenStudy (anonymous):

how about formula?

OpenStudy (compphysgeek):

the second unit that you want to change into needs to be different from the first unit. there is nothing you need to change here. If you want to change 1 ft^3 into m^3 for example that something else

OpenStudy (anonymous):